The rectangular playing area of a professional football field is 100 yards long. The perimeter of the playing area is 920 feet. The formula for the perimeter of a rectangle is P = 2l + 2w. What is the width of the playing area?
Answer by thejackal(72)

the playing area is 100 yards long. convert it to feet thus
300 feet
---------------------------------------------
|...........................................|
|...........................................|
|...........................................| W
|...........................................|
|...........................................|
|--------------------------------------------
P = 920 feet = 2l + 2W = 2(300) + 2W = 600 + 2W
(920 - 600)/2 = W